An engineering consulting office has output as service hours which has variable
cost cV & fixed cost Cf of $50 per hour &2,100,000 per year respectively. The
charge-out (selling price P) is $90 per hour & the maximum output of this firm
is 250,000 hours per year. For this firm estimate the following:
a) Is it possible to have 250,000hours yearly output?
b) Evaluate the project at full data, 10% reducing in the fixed cost, 10% reducing
in the variable cost, 10% reducing in the cost as all & 10% reducing in the unit
price.
c) At 10% interest & 20years period live estimate the present (current) worth value
of that office.
1
Expert's answer
2020-05-21T10:00:42-0400

a)"BEP=\frac{FC}{P-VC}=\frac{2 100 000}{90-50}=52 500"

"Profit=R-C=90*\times 250 000-(2 100 000+50\times250 000)=22 500 000-14 600 000=7 900 000"

above the breakeven point, so it can

b) "FC=2 100 000-0.1\times2 100 000=1890 000"

"VC=50\times250 000-0.1\times50\times250 000=11 250 000"

"P=90-0.1\times90=81"

"Profit=R-C=81*\times 250 000-(1 890 000+11 250 000)=7 110 000"

c)